11) National Auto Loan has customers throughout the 48 contiguous states. Its current policy is to have all customers mail their payments to National Auto's corporate office located in Lansing, Michigan. It is considering a regional lock-box system to collect loan payments from its customers. Changing to a regional lock-box system will decrease the company's float time from six days (on average) to three days. The company's average daily collection is $1,750,000. The bank that operates the regional lock-box system requires compensating balances totaling $800,000. Annual fees and expenses associated with the lock-box system are $350,000.

Required:

b. Compute the percentage rate of net funds freed.
Answer #1
Cash freed = (6-3)*1750000 = $5,250,000
Net cash freed = 5250000-800000 =

$4,450,000

% rate of net funds freed = 350000/4450000 = 7.87%